40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Mar 21 15:13:44 2017, gautam, Update, IMC, IMC input beam mode matching IMG_7220.JPGIMC_ModeMatch.pdfsingleLensSensitivity.pdfsensitivity.pdfIMCmodeMatch.m
    Reply  Tue Mar 21 21:59:48 2017, gautam, Update, IMC, IMC input beam mode matching IMG_3654.JPGIMG_3655.JPG
       Reply  Wed Mar 22 00:33:00 2017, gautam, Update, IMC, IMC length offset nulling offsetInvestigation.pdfoffset_summing_amp.pdf
          Reply  Wed Mar 22 16:58:25 2017, gautam, Update, IMC, WFS sensing matrix measurements 
             Reply  Thu Mar 23 01:44:53 2017, gautam, Update, IMC, WFS sensing matrix measurements IMC_WFS_segment_TF.pdfIMC_WFS_channels_TF.pdfTFsummary.pdfIMC_WFS_170322.xlsx.zip
                Reply  Thu Mar 23 08:43:11 2017, rana, Update, IMC, WFS sensing matrix measurements 
                   Reply  Thu Mar 23 23:38:58 2017, gautam, Update, IMC, MC SUS damping gains stepped down 
                      Reply  Fri Mar 24 11:26:57 2017, gautam, Update, IMC, MC SUS damping gains restored 
                         Reply  Fri Mar 24 19:04:18 2017, gautam, Update, IMC, Seismic feedforward and WFS MCL_comparison.pdfseis_comparison.pdf
Message ID: 12903     Entry time: Thu Mar 23 23:38:58 2017     In reply to: 12902     Reply to this: 12904
Author: gautam 
Type: Update 
Category: IMC 
Subject: MC SUS damping gains stepped down 

I've reduced the gains of the damping on all 3 MC SUS by a factor of 3 for overnight observation as part of the ongoing feedforward noise cancellation investigations. I will return them to the nominal values tomorrow morning.

ELOG V3.1.3-